OpenStudy (anonymous):

solve for x x^3-3x-2=0

OpenStudy (anonymous):

ok so this is a cubic which means that there are 3 x values. to find the first one you need to do trial and error. What you do is substitute random values for x until you get the left side of the equation equal to 0

OpenStudy (anonymous):

I think -1 would work

OpenStudy (anonymous):

like this, x^3 -3x - 2 = 0 (0)^3 - 3(0) - 2 = 0 0 - 0 - 2 = 0 -2 = 0 Therefore 0 isnt one of the three values of x. so now you try, plug in any value for x (id try low numbers like 1 and 2)

OpenStudy (anonymous):

(-1)^3 -3(-1) - 2 = 0 -1 +3 - 2 = 0 3 - 3 = 0 0 = 0 yep! it's a value!

OpenStudy (anonymous):

ok great! What would I do from here?

OpenStudy (anonymous):

therefore when you factorize x^3 - 3x - 2 you'll get (x + 1)( x )(x ) so now you have to divide x+1 into the cudic equation and you'll end up with a squared equation looking like this x^2 + x + c

OpenStudy (anonymous):



OpenStudy (anonymous):

do you know how to do this?

OpenStudy (anonymous):

I think I can figure it out

OpenStudy (anonymous):

ok well try it and let me know wat you get :)

OpenStudy (anonymous):

@HawkCrimson I got \[x^2 -x-2\] is that correct?

OpenStudy (anonymous):

great! now all that's left is to factorize that equation :) and get the values of x from it

OpenStudy (anonymous):

okay, I did it and I got (x-2)(x+1)

OpenStudy (anonymous):

nice so wat are your 3 values of x?

OpenStudy (anonymous):

x= -1,2,-1

OpenStudy (anonymous):

@HawkCrimson I want to thank you so much for walking me through the process I really appreciate it ! :) Thank You

OpenStudy (anonymous):

np! :)